OpenStudy (anonymous):

what does a equal in the problem 7/8a = 42?

OpenStudy (anonymous):

yes

OpenStudy (shane_b):

So is it\[\frac{7}{8}a=42\] or \[\frac{7}{8a}=42\]

OpenStudy (anonymous):

Multiply both sides by the reciprocal which is 8/7

OpenStudy (anonymous):

its the top equation. \[\frac{ 7 }{ 8 }a = 42\]

OpenStudy (shane_b):

The do as @fiedster23 said. Multiply both sides by 8/7 so that you end up with a isolated to one side.

OpenStudy (shane_b):

*The = then :P

OpenStudy (anonymous):

thank you both. i got 48 for what a equals. is that right?

OpenStudy (shane_b):

Yep

OpenStudy (anonymous):

okay thanks!!!!

OpenStudy (shane_b):

yw